If X and Y are independent Poisson variates such that
P(X =1) = P(X = 2) and P(Y = 2) = P(Y = 3) .
Find the variance of X− 2Y.
1
Expert's answer
2020-04-16T19:14:38-0400
"P(X=x)={e^{-\lambda_1}\lambda_1^x \over x!}"

Given that "P(X=1)=P(X=2)." Then


"{e^{-\lambda_1}\lambda_1^1 \over 1!}={e^{-\lambda_1}\lambda_1^2 \over 2!}, \lambda_1>0"

"\lambda_1=2"

"P(Y=y)={e^{-\lambda_2}\lambda_2^y \over y!}"

Given that "P(Y=2)=P(Y=3)." Then


"{e^{-\lambda_2}\lambda_2^2 \over 2!}={e^{-\lambda_2}\lambda_2^3 \over 3!}, \lambda_2>0"

"\lambda_2=3"

For a Poisson random variable "Z,\ Var(Z)=\lambda." Then


"Var(X)=\lambda_1=2, Var(Y)=\lambda_2=3"

Hence


"Var(X-2Y)=Var(X)+2^2Var(Y)="

"=2+4(3)=14"

Therefore

"Var(X-2Y)=14"